OpenStudy (anonymous):

Write the first ten terms of a sequence whose first term is -10 and whose common difference is -2.

OpenStudy (anonymous):

................

OpenStudy (texaschic101):

-10 -2 = -12 -12 - 2 = -14 -14 - 2 = -16 -16 - 2 = -18 -18 - 2 = -20 -20 - 2 = -22 etc....